OpenStudy (anonymous):

Find the equation of the tangent line to the curve y =√x at x = 9. (that's a square root symbol over the x)

OpenStudy (baldymcgee6):

Find the derivative and evaluate at x = 9

OpenStudy (anonymous):

the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\) replace \(x\) by 9

OpenStudy (baldymcgee6):

if it helps, think of sqrt(x) as x^(1/2) and apply the power rule to find the derivative.

OpenStudy (dape):

That's gives the slope of the tangent line, the tangent line should touch the curve at x=9, so we get: \[ y(9)=√9=a9+b \Rightarrow b=3-9a \] Where \(a=f'(9)\) is the slope and \(b\) is a constant, the line is therefore: \[ y=ax+b=f'(9)x+2-9f'(9) \] Now just find f'(9).

OpenStudy (anonymous):

i don't

OpenStudy (anonymous):

the derivative of \(\sqrt{x}\) is always \(\frac{1}{2\sqrt{x}}\) it never changes just like \(8\times 7=56\) just know it

OpenStudy (anonymous):

also \(\frac{1}{2}x^{-\frac{1}{2}}\) is useless if you want to actually compute anything

OpenStudy (baldymcgee6):

@satellite73 you would obviously simplify it further to \[\frac{ 1 }{ 2\sqrt(x) }\]

OpenStudy (baldymcgee6):

it is good to know how to get to the solution without "just knowing it".. in time, it will become a habit though, like you said. :)

OpenStudy (anonymous):

you get the solution by computing once never again might as well use \[\lim_{h\to 0}\frac{\sqrt{x+h}-\sqrt{x}}{h}\] get \(\frac{1}{2\sqrt{x}}\) and never look back

OpenStudy (baldymcgee6):

it's silly to think that one will memorize all the derivatives.

OpenStudy (baldymcgee6):

@satellite73 “I don't need to know everything, I just need to know where to find it, when I need it” ― Albert Einstein